there are 35 students in a class and the ratio of boys to girls is 3:4. How many boys and girls are there in the class?

Answer by checkley77(12844)

3:4 total=35.
3n+4n=35
7n=35
n=35/7
n=5 this is the multiplier.
3*5=15 boys
4*5=20 girls.
Proof
15+20=35
35=35